### v1

# Hesitation and Tolerance in Recommender Systems

Abstract.
User interactions in recommender systems are inherently complex, often involving behaviors that go beyond simple acceptance or rejection. One particularly common behavior is hesitation, where users deliberate over recommended items, signaling uncertainty. Our large-scale surveys, with 6,644 and 3,864 responses respectively, confirm that hesitation is not only widespread but also has a profound impact on user experiences. When users spend additional time engaging with content they are ultimately uninterested in, this can lead to negative emotions, a phenomenon we term as tolerance. The surveys reveal that such tolerance behaviors often arise after hesitation and can erode trust, satisfaction, and long-term loyalty to the platform. For instance, a click might reflect a need for more information rather than genuine interest, and prolonged exposure to unsuitable content amplifies frustration. This misalignment between user intent and system interpretation introduces noise into recommendation training, resulting in suggestions that increase uncertainty and disengagement. To address these issues, we identified signals indicative of tolerance behavior and analyzed datasets from both e-commerce and short-video platforms. The analysis shows a strong correlation between increased tolerance behavior and decreased user activity. We integrated these insights into the training process of a recommender system for a major short-video platform. Results from four independent online A/B experiments demonstrated significant improvements in user retention, achieved with minimal additional computational costs. These findings underscore the importance of recognizing hesitation as a ubiquitous user behavior and addressing tolerance to enhance satisfaction, build trust, and sustain long-term engagement in recommender systems.
Recommendation System, User Behavior Analysis, User Retention
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 11ccs: Information systems Recommender systems
1. Introduction
In recent years, the boundaries of Internet applications have been disappearing. A notable cross-domain effect has emerged: the e-commerce platforms increasingly use short videos to promote products and boost consumer spending (Lei et al., 2021; Guo et al., 2022), while social media and video live-streaming platforms have begun engaging in product sales (Lin et al., 2019; Ladhari et al., 2020; Min Zhang and Luo, 2020; Chen et al., 2019b), which have become major revenue sources. This evolution signifies a profound transformation in traditional business competition strategies. The margins of profit opportunities for these platforms are becoming increasingly blurred, shifting the competition back to the essence of the internet: attracting more users, retaining them longer, and keeping them engaged. Platforms that can achieve this holds a greater competitive advantage and business potential (Du et al., 2019; Legner et al., 2019; Yang et al., 2023; Altman and Jiménez, 2019). Consequently, the emphasis on user retention rates has become a consensus among all practitioners (Jannach and Jugovac, 2019; Gomez-Uribe and Hunt, 2016; Wang et al., 2022).
Recommender systems significantly enhance user satisfaction and retention by providing information or products that fits user’s need (Altman and Jiménez, 2019; Jannach and Jugovac, 2019). Platforms such as TikTok, Kuaishou, and Reels (Zannettou et al., 2024; Chang et al., 2023a; Molem et al., 2024) have garnered millions of users and achieved tremendous commercial success by developing exceptional recommender systems focused on user retention. At the same time, traditional industries such as online retail (Chen et al., 2019a; Leino and Räihä, 2007), news reading (Xiao et al., 2022; Zhang et al., 2023b), and music streaming (Feng et al., 2021; Briand et al., 2021) have also introduced recommendation technologies and adjusted their business models. Consequently, the goal of retaining more users i.e., user retention, and maintaining their activity in recommender systems has received unprecedented attention, with these systems being widely used to enhance user stickiness (Wu et al., 2017; Zou et al., 2019; Cai et al., 2023; Xue et al., 2023b; Silveira et al., 2017; Zhang et al., 2021b; Ding et al., 2023a; Gomez-Uribe and Hunt, 2016).
Conventionally, recommender systems classify user responses to recommended content as either interested or disinterested, treating these responses as positive and negative samples during optimization. Because click behaviors are interpreted as signals of interest, Click-Through Rate (CTR) becomes a crucial metric for assessing recommendation accuracy (Zhang et al., 2023a, 2014; Wang et al., 2015). However, the complexity and diversity of user intentions, and even emotions, complicate the correlation between intentions and the limited set of behavior types (Xia et al., 2021). For instance, a click may not only reflect interest but could also indicate a desire for more information before making a judgment. The final evaluation after clicking might still be disinterest, and the user’s attitude during the session may not clearly fall into either the interested or the disinterested category.
As a result, training recommendation models to interpret clicks primarily as positive signals can introduce uncertainty into the system. This issue is closely tied to the interaction design of the platform, especially regarding the amount and type of information exposed to users. Therefore, it is crucial to understand how this phenomenon affects user experience in recommender systems, and whether such sessions represent an overlooked state of user acceptance, which could directly influence long-term user experience, as reflected in user activity and retention rates.
To the best of our knowledge, there is no in-depth research that has thoroughly investigated this phenomenon. One possible reason for this gap is that academia often lacks corresponding signal verification, while industry, due to commercial competition, tends not to disclose too much on technical indicators. The impact of hesitation remains unclear. To address this gap, we conduct a large-scale user study, combined with dataset analysis and online A/B experiments, to examine the effects of indecision and related emotions on recommender systems.
In this study, we aim to address three research questions (RQs): (1) Does an ambiguous state exist between interest and disinterest? (2) If so, how can this state be accurately identified or captured? (3) If captured, can recommender benefit from such behavioral signals to improve retention rates?
To address RQ1, we conducted two comprehensive user surveys, gathering 6,644 and 3,864 responses, respectively. Our findings confirm that: (i) Users indeed experience a state between interest and disinterest, which we term hesitation; (ii) This hesitation, when associated with unproductive costs and decisions of disinterest, can develop into tolerance; (iii) Tolerance may diminish user enthusiasm for the platform and reduce their activities, then potentially lead to platform abandonment. Based on the processed survey data, we conclude the characteristics of tolerance signals and their application in optimizing recommender systems, which partially answered RQ2. The answer to RQ3 is based on offline analysis of existing datasets of recommender systems from e-commerce and short-video sharing platforms. The analysis concludes that an increase in tolerant behavior decreases user activity in recommender systems. We then deployed online A/B tests on a real-world video-sharing platform with millions of users to verify whether incorporating tolerance signals could reduce platform abandonment. The online results show that even low-cost strategies targeting these signals can significantly enhance next-day user retention. Training models to incorporate tolerance signals not only boosts user activity but also improves retention rates, addressing key challenges in industrial recommender systems in a cost-effective manner.
2. Related Work
Our research is related to user behavior modeling and optimization objectives in recommender systems. User behaviors, such as clicks, views, and purchases, serve as essential data for the recommender system to predict user preferences and satisfaction. Leveraging user behaviors, various optimization objectives have been proposed, i.e., predictions of click-through rate, click conversion rate, dwell time, and even long-term outcomes such as user retention. In this section, we briefly discuss existing research related to user behavior modeling and optimization objectives in recommender systems.
2.1. User Behaviors Modelling in Recommender System
Among various ways to implement and refine recommender systems, user behavior is crucial for training and evaluating recommendation models (Wu et al., 2022; Chen et al., 2023; Wang et al., 2021a). Both academia and industry concur on the key behaviors that signal interest or disinterest (Wang et al., 2023; Frolov and Oseledets, 2016; Park and Lee, 2022; Zhao et al., 2018; Liu et al., 2010). However, there remains a notable gap in the study of more nuanced, long-tail behaviors, with even less attention given to the deeper emotional interpretations of user actions (Tkalčič, 2018; Xia et al., 2021).
Preliminary work by Tkalčič (2018) provides a theoretical overview and practical insights into the impact of emotions and personality on recommender systems, concluding that user emotions significantly influence the generation and optimization of recommendations. Recently, a few studies (Pan et al., 2023; Zannettou et al., 2024; Li et al., 2022; Curmei et al., 2022) delve into behavioral interpretation and modeling through exploring the relationships among user interests (Zannettou et al., 2024), activities (Pan et al., 2023; Li et al., 2022), dwell time (Zhan et al., 2022; Zhao et al., 2024), etc., and developed various optimization objectives (Curmei et al., 2022; Li et al., 2022). Specifically, some studies propose to optimize user behavior modeling by focusing on mitigating the systematic and algorithmic biases in recommender systems (Chen et al., 2023). The systematic biases are usually suppressed by promoting equity of attention for items or improving model performances for subgroups (Beutel et al., 2019; Biega et al., 2018; Mehrotra et al., 2018; Zehlike and Castillo, 2020). The algorithm biases are commonly solved from the perspective of causal reasoning (Pearl, 2012). These solutions are roughly categorized into causal embedding (Bonner and Vasile, 2018; Guo et al., 2019; Zhang et al., 2023b), inverse propensity weighting (Christakopoulou et al., 2020; Kool et al., 2020; Bietti et al., 2021), and causal intervention (Wang et al., 2021b; Yang et al., 2021; Zhang et al., 2021a) based methods.
In addition, there are studies exploring to mine more useful user features (Cho et al., 2019; Park et al., 2017; Xiao et al., 2022) or devise multi-task frameworks (Chang et al., 2023b; Ma et al., 2018b; Tang et al., 2020; Zhao et al., 2019) to improve the personalization of recommendations and enhance user satisfaction.
Despite the valuable insights these studies provide, they often concentrate on specific optimization objectives, leaving a gap in understanding user behavioral dynamics throughout the decision-making process. To address this, our research is the first to introduce and formally define two critical concepts, i.e., hesitation and tolerance signals, within the context of recommender systems. These signals are applied in practice to capture subtle user behaviors that have been largely overlooked. Our work fills a gap in both academic and industry research by offering foundational insights into user decision-making patterns that have not yet been systematically explored.
2.2. Optimization Objective in Recommender System
Click is one of the key user behaviors in recommender system. The conversion of a click can lead to an item purchase or an advertisement view on an online platform, ultimately generating business revenue. Accurately predicting whether a user will click on an item in a given context is crucial for optimizing how items are displayed to maximize business outcomes (Zhou et al., 2018, 2019; Lin et al., 2023). This task, known as click-through rate (CTR) prediction, has been extensively studied, with numerous models proposed over the years.
Conventional CTR models, such as Factorization Machines (FM) (Rendle, 2010), model pairwise feature interactions for binary classification of clicks and non-clicks. Other approaches (Guo et al., 2017; Lian et al., 2018) extend these models by incorporating higher-order feature interactions through the use of deep neural networks (DNNs). Zhou et al. (2018) further enhance CTR models by introducing a local activation unit to capture the relevance between candidate ads and users’ historical behaviors. In DIEN (Zhou et al., 2019), the authors propose leveraging auxiliary loss in CTR models to supervise the learning of user behavior sequences at each time step. Note that, traditional CTR prediction methods typically treat all clicks equally, without distinguishing between the quality of clicks. Nevertheless, studies (Xie et al., 2023; Wen et al., 2019) also highlight that while clicks reflect users’ initial impressions, they do not necessarily indicate user satisfaction with items after further exploration. These studies suggest that post-click behaviors should be considered in CTR models to better represent user preferences.
In addition to predicting click-through rate (CTR), click conversion rate (CVR) prediction is also considered a crucial task in recommender systems. Generally, click conversion relates to the “click-to-purchase” path, which directly impacts final revenue. However, predicting CVR is fundamentally more challenging than predicting CTR due to issues such as selection bias and the sparsity of positive samples in the data. To address these challenges, Ma et al. (2018a) propose a multi-task learning approach that simultaneously trains CTR and CVR models, allowing the CVR prediction task to benefit from the abundant positive samples available in the CTR prediction task. Similarly, the authors of (Wen et al., 2020) introduce a multi-task model based on conditional probability rules to enhance CVR learning. Wen et al. (2021) further propose to decompose user behaviors into detailed actions, such as clicking on item images and engaging in conversations with sellers, providing richer signals for CVR prediction.
When it comes to video recommendation scenario, beyond clicks, watch-time prediction is regarded as a more representative metric of user engagement (Wu et al., 2018; de Sá et al., 2021; Covington et al., 2016). Covington et al. (2016) presents an industry-standard solution to watch-time prediction, where the regression problem is transformed into the weighted logistic regression. Zhan et al. (2022) further analyzes the duration bias in watch-time prediction via a causal graph and duration-deconfounded quantile-based framework. Zhao et al. (2024) introduces the counterfactual watch time concept to measure the potential duration bias and devise a corresponding counterfactual watch model to solve this problem.
Our work fundamentally differs from the aforementioned papers. Rather than focusing on immediate feedback like clicks and dwell time, we examine users’ decision-making processes, particularly their emotional responses to items after clicking. Building on this, we explore a long-term business metric: user retention. In the following section, we review existing research on optimizing user retention in recommender systems.
2.3. User Retention Optimization
User retention is typically defined as the ratio of users who revisit a system after their initial interaction (Cai et al., 2023). It has become an increasingly critical metric in the evaluation and optimization of recommender systems (Wu et al., 2017; Zou et al., 2019; Cai et al., 2023; Xue et al., 2023b; Silveira et al., 2017; Zhang et al., 2021b; Ding et al., 2023a; Gomez-Uribe and Hunt, 2016). Unlike immediate feedback mechanisms, such as point-wise (He et al., 2021; Zamani and Croft, 2020) and list-wise (Cui et al., 2021; Zheng et al., 2020) signals, which reflect user responses after a single interaction, retention captures long-term feedback resulting from multiple interactions over time (Cai et al., 2023).
Due to the nature of this delayed feedback, reinforcement learning (RL) methods are particularly well-suited to optimize user retention by addressing long-term rewards and user engagement (Zhang et al., 2022; Xue et al., 2023a; Zou et al., 2019). One of the representative works is RLUR (Cai et al., 2023), designed to address challenges such as uncertainty, bias, and delayed feedback, thereby improving retention by focusing on long-term positive outcomes. In addition, some studies have explored optimizing retention through RL by modeling mid- and long-term user behaviors (Wu et al., 2017; Zou et al., 2019; Xue et al., 2023b), investigating interpretable aspects of retention (Ding et al., 2023b), or integrating exploratory behaviors with reinforcement learning to enhance user experiences (Chen et al., 2021). Recent efforts (Kohavi and Chen, 2024) also analyze the impact of false positives in online A/B tests for recommender systems, where the mined false positives can be regarded as the retention signal.
3. User Behavior understanding through Surveys
In recommender systems, user preferences are often inferred from interactions with items, enabling refinement of the recommendation model. However, this inference process can introduce undesirable biases and errors. For instance, clicking on an item does not always indicate genuine interest; it may simply reflect a desire for more information. When users explore item details only to realize a lack of interest, interpreting the initial click as a preference signal leads to inaccurate assumptions. Consequently, such actions produce misleading feedback, potentially increasing uncertainty in future recommendations.
In addition, the decision-making process among users, based on their stated preferences of “interested” or “disinterested”, involves a series of complex and varied emotional processes. Before solidifying into definitive attitudes of interest and disinterest, individuals may incur certain costs, including time and potentially financial resources. For example, consider a scenario where a platform recommends a video that gradually loses the user’s interest until they eventually dismiss it. This scenario contrasts with a situation where the user immediately expresses disinterest and swipes away the recommended video. Determining which scenario yields a superior user experience is challenging. While the former scenario incurs a greater time cost to the user, the latter facilitates rapid disengagement based on disinterest. Consequently, there is a need to emphasize the investigation of emotional dynamics during user decision-making. This involves exploring potential novel emotional factors and their consequential influence on recommender systems.
3.1. User Survey Hypothesis
Based on the above discussion, we propose the following 4 hypotheses for our research objectives, and we verify them through user survey.
  • In the utilization of recommender systems, there exists an intermediary decision-making state, described as “hesitation”. Hesitation occurs before users can decide whether they are interested or not, and takes more costs than immediate decision.
  • Hesitation constitutes a state of the process, yet the incurred cost of hesitation may affect a user’s level of interest or disinterest in an item.
  • Hesitation is associated with increased decision-making costs. When users, following a period of hesitation, conclude that they are disinterested, this transition often gives rise to an emotion termed as “tolerance”. This emotional state is characterized by a more negative user experience compared to immediate disinterest.
  • The accumulation of tolerance results in a diminution of user trust, consequently diminishing interest in the recommender system and leading to a decrease in user activity and ultimately resulting in user attrition.
Figure 1. The potential relationship between users’ trust in the platform and their interest in its recommended items.
These hypotheses are expected to shed light on the intricate process of user emotional responses during interactions with recommender systems, potentially uncovering subtle factors that could impact the systems’ efficiency in retaining users. An illustration of our hypotheses is provided in Figure 1. By comprehending and addressing these intermediate states, particularly hesitation, and tolerance, there is potential to optimize recommender systems to better align with user requirements, and enhance user trust and retention rates within the platform.
3.2. Survey Design Principles
We conduct user survey to test the above hypotheses. It is crucial to meticulously design the survey methodology (Krosnick, 2018; Schwarz, 1999). Notably, a well-designed survey should strategically avoid many direct questions about hesitation, as they often lead to inaccurate and superficial responses, and incorrect survey results may occur due to the inclusion of erroneous options. Also, although we introduce the concept of “tolerance” in this study, we cannot use this term directly in surveys. Instead, we aim to capture tolerant behavior by describing specific scenarios and combinations of user actions, such as “assuming interest but realizing not, after further exploration” to examine its impact. Further, to avoid potential biases, we conduct two separate user surveys. Each survey is divided into five distinct sections of questions, with each section designed to sequentially capture users’ thought processes, ensuring the reliability and validity of the results. Additionally, the composition of each question set aligns with the core idea behind the questionnaire design.
In the first user survey, we collect common user behaviors that express interest and disinterest, as well as behaviors where users initially think they are interested but later realize they are not. We also assess users’ attitudes by asking them to judge the extent to which they agreed with each of our predefined answers, thereby verifying the reasonableness of our designed options. At the same time, we also investigate whether users have experienced hesitation and explore the reasons they believe contributed to it. Additionally, we outline a series of scenarios where users realize they are actually not interested after hesitating. Users are asked to evaluate these scenarios based on their own experiences and choose the ones that have impacted their use of recommender systems.
In the second user survey, we design the first three sections of questions with the same objectives as the first survey to cross-verify the validity of the conclusions drawn from the two user surveys. We then introduce two new sections of questions that focus on the impact of hesitation and tolerance, such as whether hesitation affects the level of user interest, whether increase in tolerance behaviors decreases user engagement, and ultimately leads to user attrition.
Note that the two user surveys are conducted independently at different time periods (detailed in Section 3.5). We do not intentionally ask the same group of users who completed the first survey to answer the second one. In each survey, we include a set of basic questions to understand the demographic profile of the participants, such as age, occupation, and the highest level of education. This demographic information helps us assess the distribution of both surveys, determine whether the traffic on the survey platform is healthy, and evaluate the validity of the conclusions drawn.
3.3. Design of the First User Survey
The first user survey is structured into five sections, each serving a distinct purpose as described below. The survey questions are listed in Appendix A.1.
User profile. As a required option in the survey questionnaire, we collect user’s basic information such as age, education, and occupation in this section to get a general understanding of our respondents and to evaluate the quality of the survey platform where the questionnaire is posted.
Context Introduction. This section initiates by posing elementary inquiries aimed at engaging the participants within the context of utilizing recommender systems. Participants are prompted to reflect upon their experiences with recommender systems, including aspects they found favorable and areas they identified as lacking. These inquiries are instrumental in elucidating the perceived benefits of recommender systems to users. They constitute a foundational element for the subsequent survey segments as values.
User Behavior Survey. We select two scenarios, online shopping and content browsing (in the context of short-video sharing), to collect users’ interactive behaviors when they are interested, not interested, or initially think they are interested but later realize they are not, in the recommended results. Additionally, we consider the potential impact of incidental behavior and trending topics on the recommendation results, aiming to verify whether these factors influence the recommendation outcomes and whether they harm the user experience.
Description and Assessment of the “Hesitation” State. We design questions to inquire whether users have experienced a state of hesitation regarding recommendation results, and what might be the reasons behind. Through these questions, we aim to help users recall their past experiences on recommender systems, thereby concretizing the psychological behavior of “hesitation”. The answers facilitate our subsequent understanding of user decisions and corresponding psychological states following the “hesitation” phase.
Recognition of Behavioral Combinations. We create a series of hypothetical scenarios, primarily within the contexts of online shopping and content browsing, in which users encounter recommendation results that they initially believe to be of their interest but later realize are not. We ask users to identify whether these behavior combinations reflect their specific psychological states based on their own experiences. The options are designed to be simple and straightforward, with the aim of obtaining reliable results.
3.4. Design of the Second User Survey
In the second survey, the first three sections include the same questions as those in the first survey to validate responses from different user groups. Recall that the two surveys are designed to be conducted separately at different time periods. The remaining two sections of the second survey focus more on the impact of hesitation. Specifically, we have the following two sections of new questions in the second user survey (see Appendix A.2 for the list of questions).
Further Investigation into Hesitation. In this section, the questions are designed to ask whether users generally prefer items they have thought about carefully (i.e., experienced hesitation before realizing they like them), or if they trust their first instinct and prefer items that do not cause hesitation, we aim to test the fourth hypothesis in Section 3.1, which is the inverse relationship between hesitation behavior and user’s satisfaction level with the platform. Additionally, we include several scenarios related to “assuming interest but realizing a lack of it after further exploration” (i.e., tolerance, as we defined), to help respondents engage with “tolerance” scenarios in preparation for the next part of the survey. The questions are designed differently for e-commerce and content browsing applications.
Tolerance and Its Effect on Retention Rates. In this section, we ask questions about the impact of tolerance, specifically the accumulation of the phenomenon of “assuming interest but realizing a lack of it after further exploration”. A structured series of inquiries is used to evaluate how the gradual increase in tolerance toward recommended content affects the overall user experience on the platform, potentially leading users to permanently leave the platform.
We list some of the questions from both surveys in Table 1, corresponding to $Q1$ through $Q9$. Respondents can select a Yes or No answer based on their past experiences with recommender systems. The full list of questions from both surveys is available in Appendix A. Due to business reasons, only the survey results (e.g., the ‘Yes’ rate as in Table 1) for the questions included in the main content of the paper are provided.
Table 1. List of 9 questions ($Q1$$Q9$) and their responses: Users can answer “yes” or “no” to these questions. Questions $Q1$$Q3$ are from the first survey, and the remaining questions are from the second survey.
No. Question “Yes” Rate
Q1 Have you ever hesitated about information, content, or items recommended to you? (Hesitation refers to being unsure of your interest at the moment and needing additional information or time to make a final decision.) 94%
Q2 After being recommended a product, if you click to view the item but do not take further actions such as adding it to your cart or making a purchase, does this imply that you are not truly interested in the product? 90%
Q3 When presented with a piece of content, if you merely skim through it quickly or use the speed-watching feature, result in a viewing percentage much lower than your usual habits, does this suggest that you are not genuinely interested in the content? 88%
Q4 Question: In an online shopping scenario, when you are already well-informed about the basic product information, which of the following actions better reflects a higher level of interest in the product? Option: You know this is the product you want and place the order directly without much hesitation. 60%
Q5 After spending a considerable amount of time viewing recommended content or products and then realizing you are not interested, do you feel frustrated, annoyed, or that your time has been wasted? 59%
Q6 Do these negative emotions become more intense as Q5’s increases? 80%
Q7 If the scenario where “you initially assume you are interested, spend time exploring it, but then realize you are not” occurs more frequently, would it lower your evaluation of the recommendation feature? 67%
Q8 If the situation where “you initially assume you are interested, invest time to understand it, only to find out you are not” becomes more common, would it gradually lead to boredom and prompt you to exit or close the platform? 70%
Q9 If you consistently encounter recommendations that fall under the scenario of “initially assuming interest but realizing after exploration that you are not,” would you ultimately choose to stop using the platform? 70%
(a) Respondents of the first survey
(b) Respondents of the second survey
Figure 2. Demographics of respondents from both surveys include age, highest level of education (completed or in progress), and occupation. User profiles remain largely consistent across both surveys.
3.5. Survey Distribution and Collection
The user survey was distributed through CSDN [^1] https://www.csdn.net/, one of the largest developer communities in China. The first survey was launched on March 7, 2024, with the questionnaire link displayed in a banner on the CSDN homepage for 17 consecutive days. During this period, we also sent messages to all logged-in users, encouraging them to participate. Additionally, the survey was promoted through CSDN’s official account on WeChat, a widely used social media platform in China. A total of $6,644$ responses were collected. The second survey, conducted via the same platform and channels, ran for five days starting May 13, 2024, and received $3,864$ responses.
To ensure that the sample accurately represented the primary users of recommender systems, a preliminary screening question was included in both surveys. This question asked respondents whether they were aware of their engagement with recommender systems in their daily activities. We filtered out users who reported being unaware of recommender systems, resulting in $5,556$ and $2,879$ valid responses, for first and second survey respectively. We believe that the number of responses to both surveys is sufficient for reliable data analysis. Additionally, the basic demographic profiles of the users from both surveys are largely consistent, as reported in Figure 2.
4. Survey Analysis and Findings
We have collated and thoroughly analyzed the results of the two user surveys and drawn several conclusions. We found that clicking behavior does indicate a certain degree of user interest, but it ultimately depends on the user’s final judgment after obtaining sufficient information on the item, which is influenced by the platform’s interaction design. For example, on an autoplay video platform, users can explore content in depth without needing to click, and this is also a key way for users to express interest. Once users have gathered sufficient information about an item (e.g., after watching a video for a while), their final attitude becomes crucial. This is because the process of obtaining information involves a certain cost in terms of time or even money. If users invest this cost and are left with meaningless information, their goodwill and trust in the platform will diminish, leading to a decrease in user activity and retention.
In the following, we detail our findings from the survey results. Recall that the first three sections of questions are the same for both surveys, and we received very similar responses across both surveys for these questions, with the differences of ‘Yes’ rate smaller than 2%. If a question appears in both surveys, we report the numbers from the first survey due to the larger number of responses. However, the second survey obtained $2,879$ valid responses, which provides a sufficient sample size for drawing conclusions. Our findings are presented from four key perspectives: 1) the value proposition of recommender system, 2) click behavior and user attitudes, 3) presence of hesitation, and 4) the definition and impact of tolerance.
Note that, since our survey questions are designed for two scenarios, online shopping and content browsing (in the context of short-video sharing), the conclusions we draw apply to these two scenarios, and may or may not be generalizable to other contexts.
4.1. The value proposition of recommender system
There are 62% of respondents favoring recommender systems for optimizing fragmented time use and maximizing the provision of relevant information. Additionally, 48% express interest in seeing refreshing content that inspires life, while 42% prioritize saving time in acquiring valuable information. Meanwhile, 56% of respondents express concerns about privacy, 49% report dissatisfaction with receiving disinterested, outdated, or repetitive content, and 43% complain that incidental behaviors have seriously affected recommendation results, leading to irrelevant or incorrect items.
Finding 1.
High-quality and efficiency are the key values of recommender systems.
Based on the results summarized above, the primary requirement for recommender systems is to meet users’ need for relevant and valuable information during fragmented time slots, even when there is no explicit intention to acquire information.
Figure 3. The relationship between users’ trust in the platform and their interest in recommended items.
4.2. Click behavior and user attitudes
For online shopping, 68% of users reported that when they see a product they are interested in, they will click into the detail page for further information. Additionally, 60% of users said they take further actions on products they are interested in, such as adding them to the shopping cart, bookmarking, or directly purchasing. In the content recommendation channel, 69% of users said they express interest by reading more content, while 64% said that further interaction with the platform (such as liking, bookmarking, and sharing with friends) indicates greater interest. Furthermore, 63% of users expressed long-term interest by following content creators. Conversely, stopping the learning process about an item was the most common way to express disinterest, with 69% of respondents agreeing. Additionally, using a video speed multiplier or progress bar was identified as a more subtle way of signaling disinterest to the system. Despite these patterns, objective data reveals that 54% of users who clicked on “item details” did not proceed further, indicating disinterest.
Finding 2.
Clicking is the entry point to learn more about an item, but the action of clicking does not always mean interest.
In short, while clicking is a common way to acquire more information, it does not necessarily indicate final interest. This highlights the importance of product interface design in accurately interpreting user behavior.
4.3. Presence of hesitation
The decision-making process requires cognitive engagement from users, and an overly prolonged period of cognitive engagement can be characterized as hesitation. In our survey, only 6% of users reported never hesitating about the information they were recommended, indicating that hesitation is a very common phenomenon. Then, 64% said that hesitation often arises when the current information fails to help them make a judgment or when they have doubts about the perspective and quality of the information. Additionally, 67% of users reported questioning the authenticity of the information upon seeing words indicating “advertisement”.
Finding 3.
Hesitation is common, and reduces satisfaction with recommender systems.
The hesitation process itself increases the user’s information acquisition cost, which contradicts the value proposition of recommender systems and affects the user experience. As a result, 60% of respondents indicated a preference for items without a hesitation process, suggesting that hesitation reduces users’ interest levels (see Table 1). An illustration of our feedback supporting this hypothesis is shown in Figure 3.
4.4. The definition and impact of tolerance
When users experience hesitation followed by a lack of interest, they often display heightened negative emotions due to the wasted time and resources. In our survey, 59% of respondents reported feeling frustrated in such situations, and 80% noted that these negative feelings intensify with repeated experiences (see Table 1). Additionally, we observe that a state of tolerance often coexists with indicators of interest but eventually transitions into signals of disinterest, as illustrated in Figure 4. More than half of respondents indicated that these behaviors initially sparked their interest, but, upon further exploration, they found the items uninteresting and felt their time had been wasted, a phenomenon we interpret as “tolerance sessions”.
Figure 4. The process of hesitation (and tolerance) on the recommended item/content.
Tolerant behaviors reduce user activity.:
From the survey results, 67% of users indicated that as the number of tolerable situations increased, their positive feelings toward the recommender system’s functionality decreased, leading them to question its effectiveness. Additionally, 70% of users reported that increased tolerance led to boredom, reduced interaction, and a likelihood of quitting the recommendation function, potentially resulting in platform abandonment.
Tolerant behaviors directly affect user retention.:
From the survey, 70% of users indicated that if tolerance continues to increase, they would leave the platform and not return. This leads us to an important conclusion of the study: the tolerance phenomenon from the user’s perspective directly impacts the platform’s retention rate.
Typical signals of tolerant behavior.:
According to the survey, 90% of users agree that in an online shopping scenario, clicking on a product without taking further actions, such as adding it to the cart, bookmarking, or placing an order, indicates a lack of genuine interest (see Table 2). Similarly, 88% of users agree that in video browsing, when a video is watched but not for a sufficient amount of time, or is viewed with actions like double-speed playback or skipping, it reflects a lack of interest (see Table 3). We interpret these behaviors as signs of diminished interest after initial engagement, adding to the user’s cognitive load. Consequently, we identify these signals as clear indicators of tolerance under the two surveyed scenarios.
Table 2. User interactions with E-commerce recommender system in different attitudes
Attitude User interactions Percentage (%)
Interest $\bullet$ Click to view the product details page, examining specifics such as style, model, and reviews. 68%
$\bullet$ After reviewing the product details, add the item to cart/favorites or proceed with a purchase. 60%
Disinterest $\bullet$ Simply ignore and swipe away the recommendation. 70%
$\bullet$ Click the “Not Interested” option to inform the system to reduce similar recommendations. 65%
Tolerance $\bullet$ Click to view the item but do not take further actions such as adding it to cart or making a purchase. 90%
Table 3. User interactions with content-aware recommender system in different attitudes
Attitude User interactions Percentage (%)
Interest $\bullet$ Click to view details and fully explore the recommended content. 69%
$\bullet$ Engage with the platform by liking, commenting, sharing with friends, or saving/bookmarking the content. 64%
$\bullet$ Feel inclined to watch more and proactively follow or subscribe to the content creator. 63%
Disinterest $\bullet$ Disregard it and quickly swipes it away. 69%
$\bullet$ Indicate disinterest to the system, requesting fewer similar recommendations. 63%
$\bullet$ Immediately switch to other channels or functions, such as using the search feature or checking trending topics. 60%
Tolerance $\bullet$ Merely skim through it quickly or use the speed-watching feature, resulting in a viewing percentage much lower than usual habits. 88%
5. Tolerance Behavior
Based on our survey findings, reducing user tolerance and improving information acquisition efficiency can enhance retention rates in recommender systems. Thus, it is crucial to minimize recommendations that encourage tolerance behavior at a systemic level.
5.1. Identification of Tolerance Behavior
The user survey shows that tolerance signals begin with signs of interest and end with disinterest behaviors, as illustrated in Figure 4. Thus, treating a user’s interaction with a single item as a discrete session, the clear shift from interest to disinterest can be used to identify the tolerance process.
Although users generally exhibit interest in similar ways, specific methodologies are needed to quantify these behaviors, particularly for qualitative descriptors of “full engagement” with an item, which are challenging to measure. As discussed, initial item checking typically signifies the onset of interest, while engrossed browsing or other positive actions indicate continued interest. Conversely, ceasing to check suggests disinterest. The challenge lies in defining appropriate positive and negative signals. For example, in video recommendations, a browsing duration can be considered a positive signal of engagement, whereas durations reflecting “superficial engagement” (e.g., fast forwarding) are treated as negative indicators. However, individual differences, such as browsing habits and patience levels, add complexity. A highly patient user might engage with up to 80% of content they find uninteresting, while an impatient user might disengage after only 20%. This variability highlights the difficulty of applying uniform standards of engagement tolerance across different type of users. Additionally, the significance of engagement duration depends on the content’s total length. For instance, five seconds is substantial for a 10-second video but negligible for a 10-minute video.
Considering these factors, this study adopts the personalized framework of the recommender system, establishing individual-specific judgment criteria. We assume that the personalities of users, particularly long-term platform participants, are relatively stable and that their interaction patterns with content remain consistent. Therefore, in the context of video recommendations, it is reasonable to infer users’ attitudes from their average browsing ratios, a conclusion supported by Zannettou et al. (2024), which validated that users’ behavioral habits are stable. Additionally, in e-commerce, users’ attitudes can be inferred based on whether they favor or add items to their cart after clicking on them. Specifically, if a user does not engage with certain content at all, it is classified as a negative sample; if the user engages but the browsing ratio is below their average for similar content or shows no further positive behavior, it is considered superficial engagement and classified as a tolerance sample; if the user engages and the viewing ratio exceeds their average or there is further positive behavior, it is considered deep engagement and classified as a positive sample.
5.2. Offline Dataset Analysis: Tolerance Behaviors
According to the survey, over $60\%$ of participants reported that an increase in tolerance would lead to a decline in interest in the content. Consequently, they are likely to spent less time on the platform and may eventually discontinue its use. Here, we further understand user behaviors through a study conducted on two offline datasets. Specifically, we examined tolerance behaviors in Taobao [^2] https://tianchi.aliyun.com/dataset/42 and Kuaishou [^3] https://www.kuairand.com datasets, which represent two distinct platforms.
5.2.1. Datasets
Taobao is an e-commerce platform offering millions of items for sale across various categories. User interactions with items on the platform typically involve behaviors such as clicking, adding to cart, or adding to favorites before making a purchase. As discussed in Section 1, clicking on an item signals exploration, but it does not necessarily reflect the user’s definitive interest in the item. On an e-commerce platform like Taobao, a purchase is a stronger indication of a user’s interest in the item. Similarly, adding an item to the cart or favorites also strongly suggests interest, as it indicates an intent to purchase. In this study, we define clicking without any subsequent actions (e.g., adding to cart, favorites, or purchasing), as negative feedback from the user. We term this behavior as “tolerance behavior”. We clarify that tolerance is fundamentally different from not clicking on a product. When a user chooses not to click, they demonstrate disinterest without investing time. In contrast, tolerance behavior occurs when a user spends time exploring a product in search of potential interest but ultimately feels that their time was wasted, potentially diminishing their overall experience.
Kuaishou is a video-sharing mobile app that allows users to watch short videos on various topics. However, like Taobao, simply clicking on or watching a video doesn’t strongly indicate a user’s interests. A better measure is the time spent watching, a longer watch time usually shows greater interest. This watch time also depends on users’ browsing habits (e.g., patient vs impatient users) and video length (e.g., a few seconds vs a few minutes). We define “tolerance behavior” as the ratio of a user’s watch time to the video’s duration. If a user watches a video with a browsing ratio below their average, it is considered tolerance behavior.
Table 4. Time periods selected for analysis and the corresponding users and items in reference week.
Dataset Reference Week Investigation Week #User #Item
Taobao $1$ $7$, Jun $8$ $14$, Jun $143,475$ $185,315$
$15$ $21$, Sep $22$ $28$, Sep $143,514$ $203,057$
Kuaishou $9$ $15$, Apr $16$ $22$, Apr $26,480$ $5,429,071$
$23$ $29$, Apr $30$ Apr - $6$ May $26,885$ $6,001,034$
5.2.2. Experimental Setting
By analyzing tolerance behaviors, we can gain valuable insights into their influence on user experience and, consequently, on retention rates and the overall business potential over time. However, due to the lack of exposure data in the Taobao and Kuaishou datasets, it is infeasible to directly observe user retention within these datasets. Therefore, we focus on user engagement rather than retention rates.
Engagement serves as a useful indicator of whether a user is likely to continue being active on the platform in the future. Specifically, for Taobao, we use the number of items clicked as a measure of engagement and analyze how the number of clicks without subsequent actions impacts user engagement. For Kuaishou, engagement is measured by the number of videos watched, and we investigate how the browsing ratio affects user engagement.
For both datasets, we select two time periods, each spanning two weeks. Each two-week period is further divided into a “reference week” and an “investigation week.” During the reference week, we observe user engagement (i.e., the number of items clicked on Taobao and the number of videos watched on Kuaishou) and tolerance levels (i.e., the number of items clicked without subsequent actions on Taobao and the average browsing ratio on Kuaishou). In the investigation week, we assess whether there is a decline in user engagement compared to the reference week. Table 4 presents the selected time periods along with the corresponding user and item statistics. [^4] It is worth noting that the selection of time periods is random, and the results remain consistent even when different time periods are chosen.
(a) Taobao
(b) Kuaishou
Figure 5. Tolerance behaviors in reference week vs proportion of users who show a decrease of engagement in investigation week.
5.2.3. Tolerance Behavior vs Engagement
In Figure 5, we categorize users based on their tolerance-related behaviors and present the proportion of users within each group who exhibit a decrease in engagement. Specifically, on Taobao, users are grouped by the number of items they clicked on without taking any subsequent action, while on Kuaishou, they are grouped by their average watch ratio during the reference week.
On Taobao, a higher number of items clicked without subsequent actions within a week correlates with increasing tolerance behaviors and a more significant reduction in user engagement. This pattern indicates that when users spend more time on items that do not result in a purchase, their overall activity on the platform tends to decline. Conversely, users who exhibit fewer than 10 tolerance behaviors in a week are more likely to remain active on the platform. Similarly, on Kuaishou, a reduction in the browsing ratio is associated with increase in user tolerance. Users are less inclined to engage with videos they find unappealing. As tolerance levels rise, a growing proportion of users demonstrate reduced activity on Kuaishou, potentially leading to user attrition.
Figure 5 also illustrates that the trend of decreased user engagement with increasing tolerance is consistently observed across different time frames. This finding suggests that tolerance levels significantly influence user satisfaction and retention rates on these platforms. Hence, although tolerance behavior is associated with exploratory actions, e.g., clicks on Taobao or video watching on Kuaishou, it does not necessarily reflect definitive interest. Instead, it provides a weaker signal of user preference on the items. Users often show initial interest and an intention to explore items at first glance but may not take further actions, such as making a purchase or continuing to watch. This indicates that only certain features of the items appeal to them, while others fail to sustain their engagement.
6. Tolerance Behavior for User Retention
The findings from the questionnaire, as presented in Section 4, emphasize a critical insight: tolerance behavior towards recommended items negatively impacts user loyalty to the platform. This underscores the importance of addressing the tolerance signal as a means to optimize recommender systems. By effectively managing this signal, platforms can enhance user retention, which is an essential business metric (Jannach and Jugovac, 2019). In this section, we detail our experimental approach to integrating tolerance behavior into the recommender systems, aiming to improve their performance.
Our initial plan is to conduct both offline and online experiments. Specifically, the offline experiment involves widely used benchmark datasets sourced from e-commerce, video, news, and movie platforms. We aim to design a method that incorporates tolerance signals as a key feature and trains models using historical data over the first $T$ days to predict user exposure on the ${(T+1)}^{th}$ day. In this setting, exposure signifies retention, while a lack of exposure denotes churn.
Unfortunately, existing offline benchmark datasets (Harper and Konstan, 2015; Gao et al., 2022) do not provide sufficient granularity for predicting whether a user will open the app on day $T+1$. Their user-item interactions are strictly retrospective. This makes it infeasible to observe how retention might change when we differentiate tolerance behaviors from other behaviors driven by interest or disinterest. Given these limitations of offline benchmark datasets, we instead focus on online experiments. Online experiments allow us to explore the intrinsic connection between tolerance signals and user retention more instantly. Meanwhile, online A/B testing in industrial recommender systems enables immediate monitoring of user responses, ensuring that we accurately measure the impact of algorithm modifications on user retention.
6.1. Instantiation of Tolerance Behavior
According to our survey findings, a user’s failure to explicitly reject an item does not sufficiently constitute a strong indication of interest or a clear positive signal. Instead, this nuanced response should be regarded as a distinct category of signals that lie outside the realm of clear positivity. Thus, we propose to classify these instances of tolerance behavior as either “weak positive” signals or even “negative” indicators of user preference. By incorporating this refined categorization, recommender systems gain the capability to interpret user feedback with heightened granularity. It enables recommendation algorithms to better capture the subtle nuances of user interactions and preferences.
Conventionally, the objective function of a recommendation model is to optimize the following binary cross-entropy loss:
(1) $\mathcal{L}=-\sum_{(u,i)\in\mathcal{Y}}y_{ui}\cdot\log\hat{y}_{ui}-\sum_{(u,j)\in\mathcal{Y}^{-}}(1-y_{uj})\cdot\log(1-\hat{y}_{uj})$
where $\mathcal{Y}$ denotes the positive user-item interaction set and $\mathcal{Y}^{-}$ denotes the negative set. Specifically, $(u,i)$ refers to a user-item pair where user $u$ has interacted with item $i$, such as by clicking on it. Conversely, $(u,j)$ indicates a user-item pair where user $u$ has not interacted with item $j$, and $y$ is the ground-truth label and $\hat{y}$ represents the prediction of recommendation model.
To consider the impact of tolerance signals, we reformulate the conventional objective of recommendation models by introducing the tolerance signals. Specifically, in the video recommendation scenario, we categorize user behavior according to the nature of their engagement. We define the positive sample as a click with follow-up active actions. For instance, click and watch beyond the user’s average viewing percentage. In contrast, a tolerance sample emerges when a user click without follow-up active actions, such as click and watch for a duration shorter than the user’s average viewing percentage. Any interaction without a click is labeled as a negative sample. In this case, we further divide the conventional positive signals into positive and tolerable signals. Hence, the positive samples are denoted as $\mathcal{Y}=\mathcal{Y}^{P}\cup\mathcal{Y}^{T}$, where $\mathcal{Y}^{P}$ indicates the clearly positive samples and $\mathcal{Y}^{T}$ indicates the tolerance samples. Through online experiments, we investigate two distinct strategies to handle the tolerance signals.
Tolerance samples as negatives. Here, we treat the tolerance sample as a negative signal and directly merge $\mathcal{Y}^{T}$ and $\mathcal{Y}^{-}$ to form the new set of negative samples. Accordingly, we modify Equation 1 to obtain the refined objective as:
(2) $\mathcal{L}=-\sum_{(u,i)\in\mathcal{Y}^{P}}y_{ui}\cdot\log\hat{y}_{ui}-\sum_{(u,j)\in\mathcal{Y}^{-}\cup\mathcal{Y}^{T}}(1-y_{uj})\cdot\log(1-\hat{y}_{uj}).$
Tolerance samples as weak positives. A tolerance signal arises directly from a user’s clicking behaviors, making it fundamentally distinct from a negative signal (i.e., non-click). However, the tolerance signal lacks the robust, follow-up engagement that characterizes a positive signal. As a result, we establish a clear precedence: “positive signal” ¿ “tolerance signal” ¿ “negative signal”. To formally incorporate this hierarchy, we employ a discounting mechanism:
(3) $\mathcal{L}_{T}=-\sum_{(u,i)\in\mathcal{Y}^{P}}y_{ui}\cdot\log\hat{y}_{ui}-\beta\times\sum_{(u,k)\in\mathcal{Y}^{T}}y_{uk}\cdot\log\hat{y}_{uk}-\sum_{(u,j)\in\mathcal{Y}^{-}}(1-y_{uj})\cdot\log(1-\hat{y}_{uj}).$
where $\beta<1.0$ serves as the discounting and down-scaling hyperparameter of tolerance signals. In our evaluation, $\beta$ is dynamically determined based on the user’s content viewing completion rate against the average user completion rate, and then normalized to the interval $[0,1]$. Notably, the primary difference between Equations 1 and 3 is that tolerance signals are separated from the original positive samples, allowing them to be treated as “weak positive samples”.
6.2. Online Experiments on User Retention
Figure 6. Experiment platform, Youku App, an online video platform with tens of millions of active users.
To study the impact of tolerance samples on user retention, we have conducted a series of experiments targeting both optimization objectives in Equations 2 and 3 respectively. The results of the experiments aim to reveal the relationship between tolerant behavior and user retention.
6.2.1. Experiment Setting
We partner with the Youku App video platform, which has tens of millions of active users, to evaluate our algorithm. The experiments are conducted on the double-column recommendation list on the second screen of the app’s homepage, as illustrated in Figure 6. We slightly refine the platform’s ranking model (SIM) (Pi et al., 2020) by adjusting the criteria for positive samples based on our findings. Previously, a simple click counted as a positive sample. Now, based on our user study, a positive sample requires a click accompanied by a viewing duration exceeding the user’s average viewing percentage. If a user clicks but watches for less time than their average, we label this as a tolerance sample. Situations with no click remain classified as negative samples.
In this study, we conduct two rounds of online A/B testing. Each round included two experimental groups with identical experimental setups but drawn from distinct user cohorts. This approach allows for thorough cross-validation of the results, reducing the risk of bias from a single experiment. In the first round (Groups 1 and 2), we apply Equation 2, treating tolerance samples as negative signals that implied user disengagement. In the second round (Groups 3 and 4), we utilize Equation 3, interpreting tolerance samples as weakly positive signals that hinted at mild interest or acceptance. Each experimental group comprised 80,000 real users, collectively generating approximately one million page views. The control groups received traffic on a similar scale.
Our primary evaluation metric is user stickiness, assessed through the Day-2 retention rate change, i.e., the percentage of Day 1 active users who remain active on Day 2. Due to potential business changes and competitor influences, conducting long-term experiments is not feasible. Thus, we limit the testing window to seven days. Beyond use retention, we also investigate the impact of tolerance signal isolation on other video recommendation evaluation metrics, such as dwell time.
6.2.2. Main Results
Table 5 summarizes the A/B test results, where $\mathcal{R}$ represents the change in retention rate, i.e., the increase in the proportion of users who remain active after being presented with recommended content. The results indicate that adjusting the labels consistently improves the Day-2 retention rate across both independent experimental sessions. This outcome underscores the potential of integrating nuanced behavioral signals into recommender systems to capture fine-grained user preferences and engagement, thereby improving the user retention. It also highlights that more sophisticated behavioral modeling increases user satisfaction and loyalty.
Recent studies have explored the user retention problem in other platforms (Zhu et al., 2024; Ding et al., 2023b; Cai et al., 2023). For instance, Cai et al. (2023) modeled user retention optimization as a Markov decision process and developed a retention critic learning, which achieves the retention improvements by around $0.053\%\sim 0.063\%$ in the online A/B test over video recommendation scenario. Ding et al. (2023b) optimized user retention from a contrastive multi-instance learning perspective, which obtains $0.65\%$ next-day user retention improvement over content recommendation scenario. Despite different online testing platforms, our strategy significantly improves average 7-day user retention with subtle model modifications or interventions, reaching $0.67\%$ and $0.36\%$ increase in two separate online tests. These results represent a significant improvement for practical business applications given their large user base, demonstrating our solution’s robustness and effectiveness.
Table 5. Results of four online A/B experiment groups, each tested over a continuous 7-day period.
Tolerance as Negative Tolerance as Weak Positive
Group 1 Group 2 Group 3 Group 4
Date $\mathcal{R}$ Date $\mathcal{R}$ Date $\mathcal{R}$ Date $\mathcal{R}$
2024-04-20 $+0.794\%$ 2024-04-28 $+0.670\%$ 2024-08-21 $+0.27\%$ 2024-09-15 $+1.62\%$
2024-04-21 $-0.120\%$ 2024-04-29 $-0.694\%$ 2024-08-22 $+0.14\%$ 2024-09-16 $+0.13\%$
2024-04-22 $+1.368\%$ 2024-04-30 $+1.695\%$ 2024-08-23 $-1.82\%$ 2024-09-17 $+1.42\%$
2024-04-23 $-0.604\%$ 2024-05-01 $+0.576\%$ 2024-08-24 $+0.45\%$ 2024-09-18 $+0.33\%$
2024-04-24 $+0.094\%$ 2024-05-02 $-0.132\%$ 2024-08-25 $-0.59\%$ 2024-09-19 $-0.88\%$
2024-04-25 $+0.991\%$ 2024-05-03 $+0.045\%$ 2024-08-26 $+0.14\%$ 2024-09-20 $+2.12\%$
2024-04-26 $+1.915\%$ 2024-05-04 $+0.417\%$ 2024-08-27 $+2.70\%$ 2024-09-21 $+0.72\%$
Average $+0.670\%$ Average $+0.360\%$ Average $+0.15\%$ Average $+0.58\%$
Table 6. Changes in the total time users stayed on the platform during the two groups of online A/B experiments.
Tolerance as Negative Tolerance as Weak Positive
Group 1 Group 2 Group 3 Group 4
Date Dwell Date Dwell Date Dwell Date Dwell
2024-04-20 $-1.92\%$ 2024-04-28 $-0.79\%$ 2024-08-21 $+2.08\%$ 2024-09-15 $-2.16\%$
2024-04-21 $-1.54\%$ 2024-04-29 $+1.19\%$ 2024-08-22 $-1.02\%$ 2024-09-16 $-2.95\%$
2024-04-22 $-3.04\%$ 2024-04-30 $-4.63\%$ 2024-08-23 $+0.00\%$ 2024-09-17 $+0.17\%$
2024-04-23 $+1.02\%$ 2024-05-01 $+1.33\%$ 2024-08-24 $-0.98\%$ 2024-09-18 $-3.32\%$
2024-04-24 $+1.37\%$ 2024-05-02 $-1.48\%$ 2024-08-25 $+0.00\%$ 2024-09-19 $-3.99\%$
2024-04-25 $-0.88\%$ 2024-05-03 $-1.81\%$ 2024-08-26 $+3.30\%$ 2024-09-20 $-0.48\%$
2024-04-26 $-3.10\%$ 2024-05-04 $-2.77\%$ 2024-08-27 $-3.19\%$ 2024-09-21 $+0.35\%$
Average $-1.18\%$ Average $-1.28\%$ Average $+0.03\%$ Average $-1.77\%$
6.2.3. Impact on Dwell Time
The observed increase in user retention suggests improved loyalty, yet the concurrent decrease in dwell time (see Table 6) reveals critical insights into engagement strategies. While traditional recommender systems often prioritize prolonged user sessions, our findings suggest that simply extending session length does not guarantee a superior experience. Instead, the shorter dwell times may reflect the removal of hesitation or indecision, which contributes little to overall satisfaction or perceived value.
This finding highlights that the quality of engagement is more important than the duration. Long sessions do not necessarily foster loyalty unless users are satisfied with the interactions. Recommender systems should focus on delivering value efficiently, prioritizing content that resonates with users over artificially inflating session length with moments of uncertainty or dissatisfaction. In practice, this means addressing subtle behavioral cues like hesitation or tolerance behaviors. By enhancing genuine user experiences and meeting their needs more effectively, platforms can build deeper loyalty and long-term engagement, paving the way for sustainable, user-focused recommendation strategies.
7. Conclusion
Our user study revealed that hesitation frequently occurs when users deliberate over whether to engage with recommended items. This hesitation often signals ambivalence, as items that provoke it are more likely to align with a decline in overall user interest. Interestingly, hesitation is often accompanied by deeper engagement, as users may explore the item further in an effort to resolve their uncertainty. However, when the item ultimately fails to meet their expectations, this extended engagement transforms into what we define as tolerance behavior. Tolerance behavior is characterized by an initial sequence of positive actions, such as clicks or extended viewing, which eventually shift to negative actions like disengagement or dissatisfaction. This progression makes tolerance behavior a clear and actionable negative signal for optimizing recommender systems.
By explicitly incorporating tolerance behavior into our optimization framework, we tested its impact on a large-scale industrial video platform. Our A/B testing demonstrated that addressing tolerance behavior led to a significant boost in user retention rates. This improvement was achieved with minimal implementation costs, i.e., simply changing the sample labels during model optimization, highlighting the practical value and scalability of this approach. Furthermore, this method provides a deeper understanding of user behavior dynamics, emphasizing the importance of moving beyond traditional engagement metrics to consider more nuanced patterns of interaction. These insights can help drive the development of smarter, more user-centric recommendation strategies that foster long-term satisfaction and loyalty.
We acknowledge that our survey design is limited two recommendation scenarios, e-commerce and content browsing or more specifically short-video sharing. Our online experiments are limited to the latter scenario only. Whether the findings generalize to more recommendation settings remains unclear and more research effert is needed.
Appendix A Survey Questions
A.1. The First User survey
A.1.1. Part 1: User Profile Survey
  • Q1.1:
    What is your gender?
  • $\bullet$ Male  $\bullet$ Female  $\bullet$ Prefer not to disclose
  • Q1.2:
    Which of the following best describes your age group?
  • $\bullet$ 20 years old or younger  $\bullet$ 21-30 years old  $\bullet$ 31-34 years old $\bullet$ 35 years old or older
  • Q1.3:
    What is the highest level of education you have completed?
  • $\bullet$ High school or below  $\bullet$ College diploma  $\bullet$ Bachelor’s degree  $\bullet$ Master’s degree  $\bullet$ PhD or higher
  • Q1.4:
    What is your current occupation?
  • $\bullet$ Student  $\bullet$ Internet practitioner  $\bullet$ Office staff  $\bullet$ Self-employed  $\bullet$ Others
A.1.2. Part 2: Contextual Introduction
  • Q1.5:
    Do you frequently utilize recommended content and related modules (e.g., product recommendations, short video recommendations, news suggestions) while using websites and mobile applications?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.6:
    Is acquiring information through intelligent recommendation channels your primary method of obtaining information when using websites and mobile applications?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.7:
    Is purchasing goods through recommendation channels your primary approach to online shopping?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.8:
    What do you perceive as the most significant value of the recommendation feature?
  • $\bullet$ It helps kill time.
    $\bullet$ It saves time that would otherwise be spent searching for information.
    $\bullet$ It provides new and interesting content, offering inspiration or suggestions for daily life.
  • Q1.9:
    What frustrates you the most about recommendation channels?
  • $\bullet$ Privacy concerns; excessive collection of personal information by the application.
    $\bullet$ Low quality of recommended content; frequent suggestions of irrelevant, outdated, or repetitive information.
    $\bullet$ Persistent recommendations based on occasional clicks or searches; feeling annoyed without an effective feedback mechanism to avoid such recommendations.
A.1.3. Part 3: User Behavior Survey
  • Q1.10:
    When you are recommended items of interest during online shopping, how do you usually respond?
  • $\bullet$ Click to view the product details page, examining specifics such as style, model, and reviews.
    $\bullet$ After reviewing the product details, add the item to my cart/favorites or proceed with a purchase.
  • Q1.11:
    When you are recommended products that are not of interest during online shopping, how do you usually respond?
  • $\bullet$ Simply ignore and swipe away the recommendation.
    $\bullet$ Click the “Not Interested” option to inform the system to reduce similar recommendations.
    $\bullet$ Consider the recommendations to be of poor quality and may not to use the app further, opting to close it.
  • Q1.12:
    When you encounter interesting recommendations while browsing content, how do you typically respond?
  • $\bullet$ Click to view the details and fully explore the recommended content.
    $\bullet$ Engage with the platform by liking, commenting, sharing with friends, or saving/bookmarking the content.
    $\bullet$ Feel inclined to watch more and proactively follow or subscribe to the content creator.
  • Q1.13:
    When you encounter uninteresting content while browsing, how do you typically respond?
  • $\bullet$ Disregard it and swipe it away.
    $\bullet$ Briefly view or read before confirming it does not match my interest, then stop.
    $\bullet$ Indicate disinterest to the system, requesting fewer similar suggestions.
    $\bullet$ Switch to other channels or functions, such as using the search feature or checking trending topics.
    $\bullet$ Close the application.
  • Q1.14:
    If you initially find recommended content interesting but later deem it unappealing upon closer examination, how do you usually respond?
  • $\bullet$ Fast-forward to finish viewing the content in a cursory manner.
    $\bullet$ Stop watching without completing the remaining content.
    $\bullet$ Indicate disinterest in hopes of reducing similar recommendations.
  • Q1.15:
    Do you believe that occasional searches influence subsequent recommendation results?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.16:
    If you view content out of curiosity or due to misleading title, then find the content not interesting, do you think this viewing action influences future recommendations?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.17:
    If you follow trending content due to social influence, despite not having a genuine interest, do you believe this affects future recommendations?
  • $\bullet$ Yes  $\bullet$ No
A.1.4. Part 4: Description and Assessment of the “Hesitation” State
  • Q1.18:
    Have you ever hesitated about information, content, or items recommended to you? (Hesitation refers to being unsure of your interest at the moment and needing additional information or time to make a final decision.) [Q1 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q1.19:
    What are your primary reasons for hesitation?
  • $\bullet$ Unable to decide based solely on the title or thumbnail image, requiring more details about the product or content.
    $\bullet$ Having doubts regarding the quality of the product or content (e.g., product reviews, authority of the content creator, presentation quality).
    $\bullet$ The content is labeled as “Advertisement” or “Marketing”.
  • Q1.20:
    When you feel disinterested in the recommendation results, does it negatively impact your feelings toward this app, to the extent that you may consider switching to other platforms?
  • $\bullet$ Yes  $\bullet$ No
A.1.5. Part 5: Recognition of Behavioral Combination
  • Q1.21:
    After being recommended a product, if you do not linger beyond the necessary time to view the content, does this indicate that you are not genuinely interested in the product?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.22:
    After being recommended a product, if you click to view the item but do not take further actions such as adding it to your cart or making a purchase, does this imply that you are not truly interested in the product? [Q2 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q1.23:
    When presented with a piece of content, if you watch it in its entirety but do not engage with it (e.g., through likes, saving, commenting) and immediately move on to the next piece of content, does this suggest that you are not genuinely interested in that content?
  • $\bullet$ Yes  $\bullet$ No
  • Q1.24:
    When presented with a piece of content, if you merely skim through it quickly or use the speed-watching feature, result in a viewing percentage much lower than your usual habits, does this suggest that you are not genuinely interested in the content? [Q3 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
A.2. The Second Survey
Note that Parts 1 -3 of the second survey are the same as those in the first survey, and therefore are omitted.
A.2.1. Part 4: Further Investigation into Hesitation
  • Q2.18:
    In an online shopping scenario, when you are already well-informed about the basic product information, which of the following actions better reflects a higher level of interest in the product? [Q4 in Table 1]
  • $\bullet$ Place order without much hesitation.
    $\bullet$ Carefully consider whether the product meets my needs, confirming my interest before placing order.
  • Q2.19:
    In the context of online shopping, please select all behaviors that reflect the scenario where initially a recommended product looks interesting to you, but upon further exploration it is not:
  • $\bullet$ View the recommended product but do not take further actions, such as adding it to favorites, adding to cart, or making a purchase.
    $\bullet$ View the recommended product, then click “Not Interested”.
    $\bullet$ View the recommended product, then left the recommendation channel for other functions or channels.
    $\bullet$ View the recommended product, then close the app or webpage afterward.
    $\bullet$ View the product, add it to favorites or cart, but do not purchase.
    $\bullet$ Others (please specify).
  • Q2.20:
    In content browsing scenarios (e.g., news, videos), please select all behaviors that reflect the scenario where a recommended piece of content looks interesting to you, but upon further exploration it is not:
  • $\bullet$ Click on the recommended content but do not finish viewing or reading.
    $\bullet$ Click on the recommended content but only skim through it.
    $\bullet$ Click on the recommended content, then click “Not Interested” upon finding it not interesting.
    $\bullet$ Click on the recommended content and, after finding it not interesting, left the recommendation channel for other channels or functions.
    $\bullet$ Click on the recommended content, upon finding it uninteresting, close the app or webpage.
    $\bullet$ Others (please specify).
A.2.2. Part 5: Exploration of “Assuming interested but realize not after further exploration” and Its Effect on Retention Rate
  • Q2.21:
    What is the primary reason for your hesitation?
  • $\bullet$ The title or thumbnail alone is insufficient to make a decision; I need time to check more details about the product or content.
    $\bullet$ I am unsure whether the product or piece of content is what I need, and need more time to consider it carefully.
    $\bullet$ I have doubts about the quality of the product or content (e.g., product reviews, authority of the content creator, layout quality).
    $\bullet$ The content is labeled as “Advertorial” or “Marketing”.
    $\bullet$ Others (please specify).
  • Q2.22:
    After spending a considerable amount of time viewing recommended content or products and then realizing that they are interesting, do you feel frustrated, annoyed, or that your time has been wasted? [Q5 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q2.23:
    Do these negative emotions become more intense as Q2.22’s increases? [Q6 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q2.24:
    If the scenario where “you initially assume you are interested, spend time exploring it, but then realize you are not” occurs more frequently, would it lower your evaluation of the recommendation feature? [Q7 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q2.25:
    If the situation where “you initially assume you are interested, invest time to understand it, only to find out you are not” becomes more common, would it gradually lead to boredom and prompt you to exit or close the platform? [Q8 in Table 1]
  • $\bullet$ Yes  $\bullet$ No
  • Q2.26:
    If you consistently encounter recommendations that fall under the scenario of “initially assuming interest but realizing after exploration that you are not,” would you ultimately choose to stop using the platform? [Q9 in Table 1]
  • $\bullet$ Yes  $\bullet$ No